OpenStudy (anonymous):

A hot air balloon is traveling vertically up- ward at a constant speed of 2.3 m/s. When it is 19 m above the ground, a package is released from the balloon. After it is released, for how long is the package in the air? The acceleration of gravity is 9.8 m/s2 . Answer in units of s

OpenStudy (anonymous):

After it is released, for how long is the package in the air? Once released, the package is in free fall with an acceleration of a = - g = - 9.8 m/s2 We know its initial velocity and initial position vi = 2.3 m/s yi = 15 m The later position of the package is given by y = yi + vi t + (1/2) a t2 y = 15 m + (2.3 m/s) t + (1/2) ( - 9.8 m/s2) t2 Now we set y = 0 and solve for t. Right? 0=15+2.3t-4.9t^2 which is -1.53 & 2 but the answer isnt correct. What am I doing wrong? Please help?

OpenStudy (anonymous):

Am I using the right formula?

OpenStudy (asnaseer):

still reading - give me a sec...

OpenStudy (asnaseer):

yi should be 19m not 15m

OpenStudy (anonymous):

therefore 2.22

OpenStudy (anonymous):

freak yea. ugh.

OpenStudy (anonymous):

i lost two points for that stupid error....

OpenStudy (anonymous):

Thank YOU!!!!

OpenStudy (asnaseer):

:) at least you learnt something - just like I did ...

OpenStudy (asnaseer):

yw